Please help me solve this equation: x^2+4x+4 by x+2
Answer by amoresroy(361)

Please help me solve this equation: x^2+4x+4 by x+2
1st step is to simplify the numerator by factoring
x^2+4x+4 by x+2 = (x+2)(x+2)/ (x+2)
(x+2) cancels out
= (x+2)/1
= x+2